OpenStudy (zohaibtarar):

Solution please? A rocket of mass 10 kg is fired vertically upwards from the surface of the Moon. The rocket propellant burns for 5 s during the launch period. If the rocket attains a maximum height of 6907 m from the Moon’s surface calculate the average thrust that is applied to the rocket during the launch. Ignore the rotational motion of the moon. The mass of the moon is m = 7.36 x 1022 kg . The radius of the moon is R = 1.74 x 106 m. The gravitation constant G = 6.7 x 10-11 N m2 kg-2 .

OpenStudy (sachintha):

First you have to find the gravitational acceleration on moon. \(\begin{array}&\;\;\sf F=\Large\frac{GMm}{(R+h)^2}&\qquad\large\rightarrow\qquad&\sf\normalsize G~\,-~gravitational~constant\\&&\sf M~-~mass~of~moon\\&&\sf m~-~mass~of~rocket\\&&\sf R~~-~radius~of~moon\\&&\sf h~~\,-~height~rocket~reached\\&&\sf g~\;\,-~gravitational~ acceleration \end{array}\\\) \(\sf mg=\Large\frac{GMm}{(R+h)^2}\\\) \(\;\;\;\sf g=\Large\frac{GM}{(R+h)^2}\)  I am not sure about the rest of the calculations. @agent0smith

OpenStudy (shamim):

welcome to open study !!!!!!!!!!

OpenStudy (mrnood):

Strictly speaking this question cannot be answered with the information given. It clearly states that 'rocket propellant burns for 5 s ' and this means that the mass of the rocket is reducing - but it does not tell you by how much it reduces. Therefore you cannot work out the potential energy of the rocket at its maximum height.